# EmoSteer-TTS: Fine-Grained and Training-Free Emotion-Controllable Text-to-Speech via Activation Steering

### v2

@@ Abstract
Text-to-speech (TTS) has shown great progress in recent years. However, most existing TTS systems offer only coarse and rigid emotion control, typically via discrete emotion labels or a carefully crafted and detailed emotional text prompt, making fine-grained emotion manipulation either inaccessible or unstable. These models also require extensive, high-quality datasets for training. To address these limitations, we propose EmoSteer-TTS, a novel training-free approach, to achieve fine-grained speech emotion control (conversion, interpolation, erasure) by activation steering. We first empirically observe that modifying a subset of the internal activations within a flow matching-based TTS model can effectively alter the emotional tone of synthesized speech. Building on this insight, we then develop a training-free and efficient algorithm, including activation extraction, emotional token searching, and inference-time steering, which can be seamlessly integrated into a wide range of pretrained models (e.g., F5-TTS, CosyVoice2, and E2-TTS). In addition, to derive effective steering vectors, we construct a curated emotional speech dataset with diverse speakers. Extensive experiments demonstrate that EmoSteer-TTS enables fine-grained, interpretable, and continuous control over speech emotion, outperforming the state-of-the-art (SOTA). To the best of our knowledge, this is the first method that achieves training-free and continuous fine-grained emotion control in TTS.
Demo pagehttps://emosteer-$t$s.pages.dev/
Introduction
@@
Text-to-speech (TTS) aims to generate natural-sounding human speech from textual input (Tan et al. 2021; Xie et al. 2025). It has been widely adopted in various domains, including voice assistants, robotics, and podcast production. Emotion-controllable TTS (EC-TTS) enhances this capability by enabling control over the emotional tone of synthesized speech, making it more expressive and engaging. Fine-grained EC-TTS takes this further by allowing precise modulation of attributes such as emotion intensity, pitch, and prosody at a local or even phoneme level. Such detailed control is vital for applications requiring nuanced expressiveness, e.g., personalized storytelling (Rong et al. 2025), empathetic human-computer interaction (Wadley et al. 2022), and precise speech editing (Peng et al. 2024).
Figure 1: Motivations of the proposed EmoSteer-TTS. (a) Existing paradigm for speech emotion control. (b) The proposed EmoSteer-TTS offers training-free, fine-grained continuous emotion control with improved interpretability.
@@
Controlling the emotional tone of synthesized speech typically requires the simultaneous manipulation of multiple characteristics, such as pitch, energy, and prosody. Independently adjusting any of these attributes often leads to undesirable artifacts. Therefore, in the literature, existing methods commonly adopt a conditional generation paradigm, including label-based methods that incorporate discrete emotion labels (Cho et al. 2025) and description-based methods that use textual emotion descriptions (Yang et al. 2025) as additional inputs to guide the speech synthesis process.
@@
Label-based EC-TTS approaches use categorical emotion labels (e.g., anger, happiness, fear) as an additional input to control the emotional expression during training and inference. For example, StyleTagging-TTS (Kim et al. 2021) uses Sentence BERT (Reimers and Gurevych 2019) to encode short phrases or keywords as emotion labels to guide the synthesis. However, such methods rely on fixed emotion labels, offering limited flexibility in control (Cong et al. 2025). Recent studies apply strength control to emotion labels. For instance, EmoSphere++ (Cho et al. 2025) converts discrete labels into the Valence-Arousal-Dominance (VAD) vector space (Mehrabian 1980), where the origin represents a neutral state. Both the type and intensity of emotion can be controlled by adjusting the direction and magnitude of the emotional vector. However, these methods rely on large emotion-labeled datasets and often struggle to generalize to unseen reference speech (Inoue et al. 2025).
@@
On the other hand, description-based EC-TTS methods use textual prompts, such as “A girl says welcome in a happy tone”, to describe the target emotion, guiding the TTS model to generate speech that aligns with the given description. For example, CosyVoice2 (Du et al. 2024) leverages textual prompts to control emotional expressiveness, enhanced via instruction fine-tuning. Similarly, EmoVoice (Yang et al. 2025) incorporates emotion descriptions into the text context to enable fine-grained emotion control. However, such methods (Guo et al. 2023; Shimizu et al. 2024; Ji et al. 2025; Li et al. 2023b) require large-scale datasets and carefully designed training procedures. Although these methods enable finer emotion manipulation, their controllability is fundamentally limited by the finite set of human language expressions, imposing an upper bound on control granularity. Moreover, they exhibit instability due to the inherent variability of textual descriptions and the stochastic nature of token sampling in the language models used for encoding.
Figure 2: Emotion steering toward sadness using neutral reference speech. Adding a sadness vector built from 50 tokens to the activations of five DiT layers (1, 6, 11, 16, 21) in F5-TTS significantly boosts the predicted sadness probability.
@@
In summary, the aforementioned methods face two key limitations, i.e., instability/poor generalization and coarse controllability, which hinder the development of fine-grained EC-TTS. The first arises from the lack of large-scale emotional speech datasets required for effective model training. The second stems from the control strategies employed in existing methods, which restrict the precision of emotion manipulation. Furthermore, the absence of exploration in emotion representations within TTS models poses challenges for researchers seeking to understand and analyze how speech emotions are encoded.
@@
To address these limitations, we present EmoSteer-TTS, a training-free approach that enables fine-grained, continuous emotion control. Specifically, we begin by analyzing the internal emotion representations of pretrained zero-shot TTS models, such as F5-TTS (Chen et al. 2025) and CosyVoice2. These models use a Diffusion Transformer (DiT) (Peebles and Xie 2023) as the backbone and employ flow matching (Lipman et al. 2023) to generate high-fidelity mel-spectrograms. As illustrated in Fig. 2, we observe that only a subset of tokens, i.e., activations, within the model significantly influences the emotional tone of the synthesized speech. Based on this insight, we propose a lightweight and efficient algorithm to extract emotionally salient tokens, e.g., those corresponding to “sad”. After identifying these tokens, we then use the difference between emotional tokens and neutral tokens to construct steering vectors for six basic emotions (Ekman 1992). These steering vectors, combined with a user-defined strength parameter, are then used to control the emotional tone of the synthesized speech.
@@
As illustrated in Fig. 1, the proposed EmoSteer-TTS enables training-free and fine-grained emotion control, offering improved interpretability over existing approaches. The main contributions of our method are three-fold:
@@
Figure 3: Overview of the proposed EmoSteer-TTS. Emotion steering vectors and steering weights are derived from pairs of neutral and emotional reference speech samples. During inference, these vectors are used to modulate the activations of the TTS model, effectively guiding it to synthesize speech that reflects the desired emotion.
Related Work
Emotion-Controllable Text-to-Speech
@@
EC-TTS systems allow users to specify speech emotions, enabling more expressive and natural-sounding voices. Label-based methods control emotion using discrete labels (Cho et al. 2025). For instance, EmoDubber (Cong et al. 2025) uses a flow-based framework with positive/negative emotion guidance and a classifier to adjust emotion intensity. HED-TTS (Inoue et al. 2025) models hierarchical emotion distributions across speech segments, allowing multi-level intensity control. Description-based methods use textual prompts to specify emotions (Shimizu et al. 2024; Li et al. 2025; Ji et al. 2024). PromptTTS (Guo et al. 2023) employs a BERT-based encoder to extract style from prompts and guide synthesis. VoxInstruct (Zhou et al. 2024) introduces semantic speech tokens and classifier-free guidance for fine-grained control from emotion descriptions. ControlSpeech (Ji et al. 2025) models emotional styles as Gaussian mixtures, aligning text and audio via KL divergence to enable zero-shot, controllable synthesis. Some zero-shot methods, e.g., MaskGCT (Wang et al. 2025b) and Vevo (Zhang et al. 2025), can also synthesize emotional speech, but they lack direct control and instead rely on reference speech. While these approaches have significantly advanced expressive speech synthesis, they typically require large-scale training.
Activation Steering
@@
Activation steering aims to directly modulate the internal activations of neural networks, providing a means to exert fine-grained control over the behavior of pretrained models. Activation steering has shown great potential in the realm of LLMs. For example, it can be used to control the behavior of LLMs, such as enhancing the truthfulness of responses (Xiao et al. 2024; Wang et al. 2025a). Researchers can identify the mapping between the activation distributions associated with false or misleading statements and those of accurate information (Rodriguez et al. 2024). Then, during the generation process, the model’s activations are steered towards the distribution representing truth, encouraging LLMs to produce more factually correct outputs (Li et al. 2023a). Activation steering can also be used to control text-to-image (T2I) diffusion models (Li et al. 2024; Nair et al. 2023). By modifying the activations of the diffusion model towards the distribution that corresponds to a particular style, e.g., impressionist or cubist, the model can generate images with the desired aesthetic qualities (Rodriguez et al. 2024; Brack et al. 2023). This not only allows for creative expression but also has practical applications in fields like graphic design, advertising, and entertainment. Inspired by these advances in controlling LLMs and T2I models, we apply activation steering to pretrained SOTA TTS models, enabling a stable and interpretable EC-TTS approach.
Method
Overview
@@
As illustrated in Fig. 3, the proposed EmoSteer-TTS approach consists of three key stages. In the first stage, we compute activation differences from pairs of neutral and emotional reference speech samples. In the second stage, we identify the top-$k$ emotion-relevant tokens (e.g., for “happy”) to construct a steering vector and its associated weight vector. In the third stage (inference), given any unseen reference speech and input text, we control the emotion of the synthesized speech by applying the steering vector with a strength parameter to modify internal activations. The proposed method is detailed in the following subsections.
Activation Extraction
@@
This activation difference is also known as the difference-in-means (Belrose et al. 2023), which can effectively extract robust feature directions. To ensure stable steering, we normalize $\mathbf{u}^{l}$ by dividing it by its L2 norm, resulting in a unit vector: $\mathbf{u}^{l}\leftarrow\frac{\mathbf{u}^{l}}{\|\mathbf{u}^{l}\|_{2}}$. The activation differences for all target layers to be steered are defined as $\mathcal{U}=\{\mathbf{u}^{l}\mid l\in\hat{\mathcal{L}}\}$, where $\hat{\mathcal{L}}\subseteq\mathcal{L}$ denotes the set of selected layers. It is worth noting that the direction of $\mathbf{u}^{l}$ indicates the trajectory of emotional change in the feature space, while its original magnitude reflects the extent of the transition between emotions.
Note that the synthesized speech may vary in length. Therefore, we interpolate the extracted activations (token sequences) to a fixed length, which is the average activation sequence length across all $M+N$ samples. As a result, each activation has the shape [avg_seq_length,hidden_dim][avg\_seq\_length,hidden\_dim].
Steering Vector Construction
@@
where $\mathbf{x}^{l}_{a}$ is the activation corresponding to a random sentence and a reference speech sample different from those used to compute $\mathbf{u}^{l}$, and $f_{\text{r}}$ is a function that renormalizes the modified activation to preserve the original L2 norm.
After the activation modification, the model synthesizes the output sample $\hat{A}_{i}$ corresponding to token $i$. We then use a pre-trained speech emotion recognition (SER) model, i.e., emotion2vec (Ma et al. 2024), to predict the probability that $\hat{A}_{i}$ corresponds to the target emotion, denoted as $P_{\text{emotion}}(\hat{A}_{i})$. By computing $P_{\text{emotion}}(\hat{A}_{i})$ for all tokens, we obtain the probability set:
@@
$\mathbf{s}^{l}\leftarrow\mathbf{s}^{l}\odot\mathbf{m},\;\mathbf{m}_{i}=\begin{cases}1,&\text{if }i\in\mathcal{I}_{\text{top-k}}\\ 0,&\text{otherwise}\end{cases},$ (5)
@@
where $\delta$ is the Softmax function: $\delta(z_{i})=e^{z_{i}}/\sum_{j=1}^{k}e^{z_{j}}$. Finally, we get the weighted steering vector $\mathbf{\hat{s}}^{l}$:
@@
which can be used to steer speech emotions. Since most elements of the weighted steering vector are zero, $\mathbf{\hat{s}}^{l}$ lies within a subspace of the TTS model’s feature space that is specifically responsible for modeling emotional tone.
To ensure the efficiency of the token searching process, we simultaneously modify all selected layers at the same token indices, which can reduce the computational complexity from $\mathcal{O}(|\hat{\mathcal{L}}|\times avg\_seq\_length)$ to $\mathcal{O}(avg\_seq\_length)$.
Fine-Grained Emotion Control
@@
Emotion Conversion and Interpolation.
As shown in Fig. 3, given the text and reference speech, we can use the steering vector $\mathbf{s}^{l}$ and steering weight $\mathbf{w}^{l}$ to modify the activations in layer $l\in\mathcal$$\hat${L}} as follows:
@@
Method Conversion Interpolation Erasure
WER($\downarrow$) S-SIM($\uparrow$) E-SIM($\uparrow$) N-MOS($\uparrow$) EI-MOS($\uparrow$) $E$-MOS($\uparrow$)
[…]
*: Training-based, #: Training-free, -: Not supported. The top three results for each metric are indicated in boldface.
Table 1: Comparison between EmoSteer-TTS and baselines, in terms of multiple objective metrics and subject metrics.
Emotion Erasure.
One may wish to synthesize new speech samples using the speaking style or timbre from the reference speech while disregarding the emotional tone. Suppose the weighted steering vector $\mathbf{\hat{s}}^{l}$ corresponds to the emotion conveyed by the reference speech, our method achieves this by subtracting the weighted steering vector $\mathbf{\hat{s}}^{l}$ from the original activation $\mathbf{x}^{l}$, multiplied by the projection of $\mathbf{\hat{s}}^{l}$ onto $\mathbf{x}^{l}$, which can be expressed as follows:
@@
Composite Control.
EmoSteer-TTS also enables composite control over the emotional tone of synthesized speech. For example, given a reference speech sample, emotion replacement can be achieved through the following operation:
$\mathbf{\hat{x}}^{l}=f_{\text{r}}(\mathbf{x}^{l}-\beta(\mathbf{\hat{s}}^{l}_{\text{emo1}}\cdot\mathbf{x}^{l})\mathbf{\hat{s}}^{l}_{\text{emo1}}+\alpha\mathbf{\hat{s}}^{l}_{\text{emo2}}),$ (10)
which replaces the emotion emo1 with emo2. In addition, we can also realize multiple emotion steering as follows:
$\mathbf{\hat{x}}^{l}=f_{\text{r}}(\mathbf{x}^{l}+\alpha_{1}\mathbf{\hat{s}}^{l}_{\text{emo1}}+\alpha_{2}\mathbf{\hat{s}}^{l}_{\text{emo2}}+...+\alpha_{E}\mathbf{\hat{s}}^{l}_{\text{emoE}}),$ (11)
@@
EmoSteer-TTS enables fine-grained, continuous emotional control and supports multiple control strategies, representing the first training-free EC-TTS approach. Appendix A in the Supplementary Materials includes code snippets for the operations described above.
Figure 4: Results of (a) emotion conversion, (b) interpolation, (c) erasure, (d) replacement, and (e) multi-emotion steering, along with an analysis of the steering dynamics in F5-TTS. (Abbreviations: Anger, Disgust, Fear, Happiness, Sadness, Surprise)
Experiment
Datasets and Models

### v3

@@ Abstract
Text-to-speech (TTS) has shown great progress in recent years. However, most existing TTS systems offer only coarse and rigid emotion control, typically via discrete emotion labels or a carefully crafted and detailed emotional text prompt, making fine-grained emotion manipulation either inaccessible or unstable. These models also require extensive, high-quality datasets for training. To address these limitations, we propose EmoSteer-TTS, a novel training-free approach, to achieve fine-grained speech emotion control (conversion, interpolation, erasure) by activation steering. We first empirically observe that modifying a subset of the internal activations within a flow matching-based TTS model can effectively alter the emotional tone of synthesized speech. Building on this insight, we then develop a training-free and efficient algorithm, including activation extraction, emotional token searching, and inference-time steering, which can be seamlessly integrated into a wide range of pretrained models (e.g., F5-TTS, CosyVoice2, and E2-TTS). In addition, to derive effective steering vectors, we construct a curated emotional speech dataset with diverse speakers. Extensive experiments demonstrate that EmoSteer-TTS enables fine-grained, interpretable, and continuous control over speech emotion, outperforming the state-of-the-art (SOTA). To the best of our knowledge, this is the first method that achieves training-free and continuous fine-grained emotion control in TTS. Demo samples are available at https://emosteer-tts-demo.pages.dev/.
1 Introduction
@@
Text-to-speech (TTS) aims to generate natural-sounding human speech from textual input (Tan et al., 2021; Xie et al., 2025). It has been widely adopted in various domains, including voice assistants, robotics, and podcast production. Emotion-controllable TTS (EC-TTS) enhances this capability by enabling control over the emotional tone of synthesized speech, making it more expressive and engaging. Fine-grained EC-TTS takes this further by allowing precise modulation of the conveyed emotion intensity in synthesized speech. Such detailed control is vital for applications requiring nuanced expressiveness, e.g., personalized storytelling (Rong et al., 2025), empathetic human-computer interaction (Wadley et al., 2022), and precise speech editing (Peng et al., 2024).
@@
Controlling the emotional tone of synthesized speech typically requires the simultaneous manipulation of multiple characteristics, such as pitch, energy, and prosody. Independently adjusting any of these attributes often leads to undesirable artifacts. Therefore, in the literature, existing methods commonly adopt a conditional generation paradigm, including label-based methods that incorporate discrete emotion labels (Cho et al., 2025) and description-based methods that use textual emotion descriptions (Yang et al., 2025) as additional inputs to guide the speech synthesis process.
@@
Label-based EC-TTS approaches use categorical labels (e.g., anger, happiness, fear) as an additional input to control the emotional expression during training and inference. For example, StyleTagging-TTS (Kim et al., 2021b) uses Sentence BERT (Reimers & Gurevych, 2019) to encode short phrases or keywords as emotion labels to guide the synthesis. However, such methods rely on fixed emotion labels, offering limited flexibility in control (Cong et al., 2025). Recent studies apply strength control to emotion labels. For instance, EmoSphere++ (Cho et al., 2025) converts discrete labels into the Valence-Arousal-Dominance (VAD) vector space (Mehrabian, 1980), where the origin represents a neutral state. Both the type and intensity of emotion can be controlled by adjusting the direction and magnitude of the emotional vector. However, these methods rely on large emotion-labeled datasets and often struggle to generalize to unseen reference speech (Inoue et al., 2025).
@@
On the other hand, description-based EC-TTS methods use textual prompts, such as “A girl says welcome in a happy tone”, to describe the target emotion, guiding the TTS model to generate speech that aligns with the given description. For example, CosyVoice2 (Du et al., 2024) leverages textual prompts to control emotional expressiveness, enhanced via instruction fine-tuning. Similarly, EmoVoice (Yang et al., 2025) incorporates emotion descriptions into the text context to enable fine-grained emotion control. However, such methods (Guo et al., 2023; Shimizu et al., 2024; Ji et al., 2025; Li et al., 2023b) require large-scale datasets and carefully designed training procedures. Although these methods enable finer emotion manipulation, their controllability is fundamentally limited by the finite set of human language expressions, imposing an upper bound on control granularity. Moreover, they exhibit instability due to the inherent variability of textual descriptions and the stochastic nature of token sampling in the language models used for encoding.
@@
In summary, existing methods have two limitations, i.e., instability/poor generalization and coarse controllability. The first arises from the lack of large-scale emotional speech datasets required for effective model training. The second stems from the control strategies employed in existing methods, which restrict the precision of emotion manipulation. Furthermore, the absence of exploration in emotion representations within TTS models poses challenges for researchers seeking to understand how speech emotions are encoded.
Figure 1: Motivations of our work. (a) Existing paradigm for speech emotion control. (b) EmoSteer-TTS offers training-free, fine-grained continuous emotion control with improved interpretability.
@@
To address these limitations, we present EmoSteer-TTS, a training-free approach that enables fine-grained, continuous emotion control, as illustrated in Fig. 1. Specifically, we begin by analyzing the internal emotion representations of pretrained zero-shot TTS models, such as F5-TTS (Chen et al., 2025) and CosyVoice2. These models use a Diffusion Transformer (DiT) (Peebles & Xie, 2023) as the backbone and employ flow matching (Lipman et al., 2023) to generate high-fidelity mel-spectrograms. As shown in Fig. 2, we observe that only a subset of tokens, i.e., activations, within the model significantly influences the emotional tone of the synthesized speech. Building on this insight, we propose a simple yet effective algorithm to extract emotionally salient tokens, such as those associated with “sad.” After identifying these tokens, we then use the difference between emotional tokens and neutral tokens to construct steering vectors for six basic emotions (Ekman, 1992). These steering vectors, combined with an adjustable strength parameter, are then used to control the synthesized emotional tone.
@@
In summary, EmoSteer-TTS enables training-free and fine-grained emotion control, offering improved interpretability over existing approaches. The contributions of our method are:
@@
Figure 2: Adding a sadness steering vector to the activations in five DiT layers (1, 6, 11, 16, 21) of F5-TTS, conditioned on neutral speech, substantially increases the predicted sadness probability.
2 Related Work
@@
Emotion-Controllable Text-to-Speech. Unlike traditional TTS systems, e.g., VITS (Kim et al., 2021a) and VALL-E (Wang et al., 2023), that produce neutral or monotone speech, EC-TTS systems allow users to specify speech emotions, enabling more expressive and natural-sounding voices. Label-based methods control emotion using discrete labels (Cho et al., 2025). For instance, EmoDubber (Cong et al., 2025) uses a flow-based framework with positive/negative emotion guidance and a classifier to adjust emotion intensity. HED-TTS (Inoue et al., 2025) models hierarchical emotion distributions across speech segments, allowing multi-level intensity control. Description-based methods use textual prompts to specify emotions (Shimizu et al., 2024; Li et al., 2025; Ji et al., 2024). PromptTTS (Guo et al., 2023) employs a BERT-based encoder to extract style from prompts and guide synthesis. VoxInstruct (Zhou et al., 2024) introduces semantic speech tokens and classifier-free guidance for fine-grained control from emotion descriptions. ControlSpeech (Ji et al., 2025) models emotional styles as Gaussian mixtures, aligning text and audio via KL divergence to enable zero-shot, controllable synthesis. Some zero-shot methods, e.g., MaskGCT (Wang et al., 2025b) and Vevo (Zhang et al., 2025), can also synthesize emotional speech, but they lack direct control and instead rely on reference speech. While these approaches have significantly advanced expressive speech synthesis, they typically require large-scale datasets and training.
@@
Activation Steering. Activation steering aims to directly modulate the internal activations of neural networks, providing a means to exert fine-grained control over the behavior of pretrained models. Activation steering has shown great potential in the realm of LLMs. For example, it can be used to control the behavior of LLMs, such as enhancing the truthfulness of responses (Xiao et al., 2024; Wang et al., 2025a). Researchers can identify the mapping between the activation distributions associated with false or misleading statements and those of accurate information (Rodriguez et al., 2024). Then, during the generation process, the model’s activations are steered towards the distribution representing truth, encouraging LLMs to produce more factually correct outputs (Li et al., 2023a). Activation steering can also be used to control text-to-image (T2I) diffusion models (Li et al., 2024; Nair et al., 2023). By modifying the activations of the diffusion model towards the distribution that corresponds to a particular style, e.g., impressionist or cubist, the model can generate images with the desired aesthetic qualities (Rodriguez et al., 2024; Brack et al., 2022). Inspired by these advances, we explore emotion representations in pretrained zero-shot TTS models and apply activation steering, offering a stable and interpretable EC-TTS method.
3 Method
Figure 3: Overview of EmoSteer-TTS. Steering vectors and steering weights are derived from pairs of neutral and emotional reference speech. During inference, these vectors are used to modulate the activations in a TTS model, guiding it to synthesize speech that reflects the desired emotion.
3.1 Overview
@@
As shown in Fig. 3, the proposed EmoSteer-TTS approach consists of three key stages. First, we compute activation differences using pairs of neutral and emotional reference speeches. Second, we identify top-$k$ emotion-relevant tokens (e.g., for “happy”) to construct a steering vector and its associated weight vector. At inference time, given any unseen reference speech and text, we control the emotion of the synthesized speech by applying the steering vector with a strength parameter to modify internal activations. The proposed method is detailed in the following subsections.
3.2 Activation Extraction
@@
This activation difference is also known as the difference-in-means (Belrose et al., 2023), which can effectively extract robust feature directions. To ensure stable steering, we normalize $\mathbf{u}^{l}$ by dividing it by its L2 norm, resulting in a unit vector: $\mathbf{u}^{l}\leftarrow\frac{\mathbf{u}^{l}}{\|\mathbf{u}^{l}\|_{2}}$. The activation differences for all target layers to be steered are defined as $\mathcal{U}=\{\mathbf{u}^{l}\mid l\in\hat{\mathcal{L}}\}$, where $\hat{\mathcal{L}}\subseteq\mathcal{L}$ denotes the set of selected layers. It is worth noting that the direction of $\mathbf{u}^{l}$ indicates the trajectory of emotional change in the feature space, while its original magnitude reflects the extent of the transition between emotions.
Synthesized speech may vary in length. Therefore, we use nearest interpolation to align the extracted activations (token sequences) to a fixed length, which is the average activation sequence length across all $M+N$ samples. As a result, each activation has the shape [avg_seq_length,hidden_dim][avg\_seq\_length,hidden\_dim].
3.3 Steering Vector Construction
@@
where $\mathbf{x}^{l}_{a}$ is the activation corresponding to a random sentence and a reference speech sample different from those used to compute $\mathbf{u}^{l}$, and $f_{\text{r}}$ is a function that renormalizes the modified activation to preserve the original L2 norm, which ensures more stable modification Gaintseva et al. (2025).
After the activation modification, the model synthesizes the output sample $\hat{A}_{i}$ corresponding to token $i$. We then use a pre-trained speech emotion recognition (SER) model, i.e., emotion2vec (Ma et al., 2024), to predict the probability that $\hat{A}_{i}$ corresponds to the target emotion, denoted as $P_{\text{emotion}}(\hat{A}_{i})$. By computing $P_{\text{emotion}}(\hat{A}_{i})$ for all tokens, we obtain the probability set:
@@
$\mathbf{s}^{l}\leftarrow\mathbf{u}^{l}\odot\mathbf{m},\;\mathbf{m}_{i}=\begin{cases}1,&\text{if }i\in\mathcal{I}_{\text{top-k}}\\ 0,&\text{otherwise}\end{cases},$ (5)
@@
where $\delta$ is the Softmax function: $\delta(z_{i})=\frac{e^{z_{i}}}{\sum_{j=1}^{k}e^{z_{j}}}$. Finally, we get the weighted steering vector $\mathbf{\hat{s}}^{l}$:
@@
which can be used to steer speech emotions. Since most elements of the weighted steering vector are zero, $\mathbf{\hat{s}}^{l}$ lies within a subspace of the TTS model’s feature space that is specifically responsible for modeling emotional tone. To ensure the efficiency of the token searching process, we simultaneously modify all selected layers at the same token indices, which can reduce the computational complexity from $\mathcal{O}(|\hat{\mathcal{L}}|\times avg\_seq\_length)$ to $\mathcal{O}(avg\_seq\_length)$.
3.4 Fine-Grained Emotion Control
@@
Emotion Conversion and Interpolation. As shown in Fig. 3, given the text and reference speech, we can use the steering vector $\mathbf{s}^{l}$ and weight $\mathbf{w}^{l}$ to modify the activations in layer $l\in\mathcal$$\hat${L}} as follows:
@@
Emotion Erasure. One may wish to synthesize new speech samples using the speaking style or timbre from the reference speech while disregarding the emotional tone. Suppose the weighted steering vector $\mathbf{\hat{s}}^{l}$ corresponds to the emotion conveyed by the reference speech, our method achieves this by subtracting the weighted steering vector $\mathbf{\hat{s}}^{l}$ from the original activation $\mathbf{x}^{l}$, multiplied by the projection of $\mathbf{\hat{s}}^{l}$ onto $\mathbf{x}^{l}$, which can be expressed as follows:
@@
Composite Control. EmoSteer-TTS also enables composite control over the emotional tone of synthesized speech. For example, given a reference speech sample, emotion replacement can be achieved through the following operation ($\mathbf{\hat{s}}^{l}_{\text{emo}_{i}}$ is the weighted steering vector of emotion “emoi”):
$\mathbf{\hat{x}}^{l}=f_{\text{r}}(\mathbf{x}^{l}-\beta(\mathbf{\hat{s}}^{l}_{\text{emo}_{1}}\cdot\mathbf{x}^{l})\mathbf{\hat{s}}^{l}_{\text{emo}_{1}}+\alpha\mathbf{\hat{s}}^{l}_{\text{emo}_{2}}),$ (10)
which replaces emotion “emo1” with “emo2”. We can also realize multiple emotion steering:
$\mathbf{\hat{x}}^{l}=f_{\text{r}}(\mathbf{x}^{l}+\alpha_{1}\mathbf{\hat{s}}^{l}_{\text{emo}_{1}}+\alpha_{2}\mathbf{\hat{s}}^{l}_{\text{emo}_{2}}+...+\alpha_{E}\mathbf{\hat{s}}^{l}_{\text{emo}_{E}}),$ (11)
@@
EmoSteer-TTS enables fine-grained, continuous emotional control and supports multiple control strategies, representing the first training-free EC-TTS approach. Appendix A provides code snippets for the operations described above.
4 Experiment
4.1 Datasets and Models